Solve the problem. You must write two equations with two variables and show steps to solve.
Anne and Nancy use a metal alloy that is 17% copper to make jewelry. How many ounces of an alloy that is 11% copper must be mixed with an alloy that is 22% copper to form 55 ounces of the desired alloy?

Found 2 solutions by ikleyn, stanbon:
Answer by ikleyn(52776)

.
I just explained you how to solve it in
https://www.algebra.com/algebra/homework/equations/Equations.faq.question.1073595.html

https://www.algebra.com/algebra/homework/equations/Equations.faq.question.1073595.html

The rest is simple arithmetic.

By the way, it is not the tutor's duty to make calculations instead of you.

It is your duty, not my.

If you don't want to work at all (as you demonstrate by repeatedly posting your problem) - it is just your deal. Not my.


Answer by stanbon(75887)

write two equations with two variables and show steps to solve.
Anne and Nancy use a metal alloy that is 17% copper to make jewelry. How many ounces of an alloy that is 11% copper must be mixed with an alloy that is 22% copper to form 55 ounces of the desired alloy?
-----
weight equation:: e + t = 55 oz
copper equation:: 0.11e + 0.22t = 0.17*55
-----
22e + 22t = 22*55
11e + 22t = 17*55
----
Subtract and solve for e::
11e = 5*55
e = 25 oz (amt. 11% alloy needed)
-------------
t = 55-25 = 30 oz (amt. 22% alloy needed)
--------------
Cheers
Stan H.
-------------
---------------------------